# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ C.1 Correctness of Algorithm 1
  • $\mathbf{Z}_{8}$ (Line 3): As proven in [48], the independence relations tested in Line 3 uniquely define partition $\mathbf{Z}_{8}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{5,7}$ (Line 4): As proven in [48], the independence relations tested in Line 4 define partition $\mathbf{Z}_{7}$ in all structures and $\mathbf{Z}_{5}$ in structures for which $|\mathbf{Z}_{1}|=0$ (i.e., no confounders, latent nor observed, exist for $X$ and $Y$). Therefore, this test is asymptotically guaranteed to return all $\mathbf{Z}_{7}$ for arbitrary structures, and all $\mathbf{Z}_{5}$ if $\mathbf{Z}_{1}$ is the empty set.
  • $\mathbf{Z}_{4}$ (Line 5): As proven in [48], the independence relations tested in Line 5 uniquely define partition $\mathbf{Z}_{4}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ (Line 8): At this stage, the only partitions remaining under consideration are $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{5}$ (assuming that $\mathbf{Z}_{1}$ is not the empty set). By definition, no $\mathbf{Z}_{5}$ will ever be adjacent to $Y$, as this partition cannot share a causal path to $Y$ that is not mediated by $X$ [34, 47, 48]. All $\mathbf{Z}_{2\notin de(Y)}$ are descended from $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and/or $\mathbf{Z}_{4}$ (i.e., any partition that can enter into $Y$), but not from $Y$ itself. Thus, all $\mathbf{Z}_{2}$ and $\mathbf{Z}_{5}$ will be eliminated by eliminating variables not adjacent to $Y$ given $X\cup\mathbf{Z}_{4}\cup\{\mathbf{Z}^{\prime}\setminus Z\}$. This test will also result in the elimination of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are not adjacent to $Y$, including proxy confounders and proxy mediators as defined in Maasch et al. [48].
  • $\mathbf{Z}_{4\in pa(Y)}$ (Line 10): Any intervening node on the causal path from $\mathbf{Z}_{4}$ to $Y$ ($\mathbf{Z}_{4}\to\cdots\to Y$) must also belong to a partition that can enter into $Y$ ($\cdots\to Y$). By definition, this is strictly $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, to obtain $d$-separation between $Y$ and any member of $\mathbf{Z}_{4}$ that is not adjacent to $Y$, a correct conditioning set will contain members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$ and $\mathbf{Z}_{4}$. It is therefore sufficient to condition all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are already known to be adjacent to $Y$, unioned with all known $\mathbf{Z}_{4}$. Any variable that is marginally dependent on $Y$ given this conditioning set is a member of $\mathbf{Z}_{4}$ that is adjacent to $Y$.
@@
  • SDC (Line 12): By definition of the SDC (Definition 2.3), $SDC=\mathbf{1}(X\in pa(Y))$. If $X$ is not a parent of $Y$, it will always be $d$-separable from $Y$ when all their confounding and mediating paths are blocked. If $X$ is a parent of $Y$, $X$ and $Y$ will be always conditionally dependent given any conditioning set. Therefore, the SDC will evaluate to 0 if and only if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$.
C.2 Proof of Theorem 3.2
@@
Maasch et al. [48] have shown previously that the marginal and conditional independence tests used to identify $\mathbf{Z}_{4}$ and $\mathbf{Z}_{8}$ are not affected by latent variables. Intuition for this follows from the fact that these tests only require knowledge of $\{X,Y\}$ and the candidate variable $Z$. In Algorithm 1, the test for $\mathbf{Z}_{5,7}$ also relies only on knowledge of $X$, $Y$, and $Z$, and is therefore correct in the presence of latent variables. As we do not require identification of $\mathbf{Z}_{5}$ itself for downstream results, this test suffices. Otherwise, latent variables could pose identifiability problems for $\mathbf{Z}_{5}$, as in prior works [48]. (A2) alone is sufficient to ensure $d$-separation between $Y$ and its non-parents at Lines 8 and 10, guaranteeing identification of the true parents of $Y$ in $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, latent variables that are not parents of $Y$ have no impact on Algorithm 1. ∎
C.3 Proof of Theorem 3.4
@@
This theorem claims that $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ is a valid adjustment set for weighted CDE identificaiton. Per the identifiability conditions of the weighted CDE (Definition 2.5), a valid adjustment set must block all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$. As shown in Pearl [59], knowledge of all mediators for $\{X,Y\}$ that are parents of $Y$ is sufficient for identifying the CDE (i.e., knowledge of every mediator in the full causal graph is unnecessary). Thus, we show that adjusting for $\mathbf{A}_{\text{{DE}}}$ blocks all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$, does not open non-causal paths by conditioning on colliders, and blocks all frontdoor paths for $\{X,Y\}$ by conditioning on all mediators that are parents of $Y$ (Definition 2.6). We use the language of causal partitions defined in [48]. We assume that all assumptions of this work are met (e.g., (A2): all parents of $Y$ are observed in $\mathbf{Z}$).
@@
    Backdoor paths for $\{X,Y\}$ are blocked. Any confounder for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{1}$, by definition of $\mathbf{Z}_{1}$ (Table 1) [48]. All backdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all confounders that are parents of $Y$, all backdoor paths for $\{X,Y\}$ will be blocked when conditioning on $\mathbf{A}_{\text{{DE}}}$.
@@
    Backdoor paths for $\{\mathbf{M},Y\}$ are blocked. Any confounder for $\mathbf{M}$ and $Y$ will be in $\mathbf{Z}_{1}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{4}$, as these are the only causal partitions with edges pointing into $Y$. Set $\mathbf{A}_{\text{{DE}}}$ by definition contains all members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are parents of $Y$. Given (A2), all backdoor paths between $\mathbf{M}$ and $Y$ are guaranteed to be blocked.
@@
    Frontdoor paths are blocked. Any mediator for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{3}$, by definition of $\mathbf{Z}_{3}$ (Table 1) [48]. All frontdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all mediators that are parents of $Y$, all frontdoor paths will be blocked.
@@
    Non-causal paths through colliders are blocked. Though $\mathbf{Z}$ can contain colliders in $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{A}_{\text{{DE}}}$ does not contain members of $\mathbf{Z}_{2}$. Thus, non-causal paths will not be opened by conditioning on $\mathbf{A}_{\text{{DE}}}$. We show in Section C.1 that Algorithm 1 is asymptotically guaranteed to remove $\mathbf{Z}_{2\notin de(Y)}$ from the final adjustment set.
Appendix D Empirical Evaluation Methods
@@
Figure D.1: Eighteen-node DAG used to evaluate the ability of LD3 to recover the true $\mathbf{A}_{\text{{DE}}}$, assess the SDC, and estimate the weighted CDE.
Baseline Evaluation
Baselines were evaluated on the basis of F1, precision, and recall with respect to the true members of $\mathbf{A}_{\text{{DE}}}$. Among baselines, only LDECC is designed to return the parents of $Y$. Additionally, both PC and LDECC return results relative to a Markov equivalence class (MEC). Therefore, we evaluate baselines both with and without post-processing to adjust for the task of $\mathbf{A}_{\text{{DE}}}$ prediction. This allows for out-of-the-box evaluation and a more charitable evaluation that is tailored for the task. In contrast, LD3 only needs to be run out-of-the-box. The two evaluation methods per algorithm are as follows: PC (PC and PC) [38], [^8] https://github.com/keiichishima/pcalg LDP (LDP and LDPpa) [48], and LDECC (LDECC and LDECC) [31]. We define the evaluation for each below.
@@
    PC: Without post-processing, we retain the parents of $Y$ in all DAGs in the MEC.
@@
    PC: With post-processing, we retain the parents of $Y$ in any DAG in the MEC.
@@
    LDECC: Without post-processing, we retain the variables labeled as parents of the target.

### v3

@@ B.1 Proof of Theorem 1 (Correctness of Algorithm 1)
  • $\mathbf{Z}_{8}$ (Line 3): As proven in (Maasch et al. 2024), the independence relations tested in Line 3 uniquely define partition $\mathbf{Z}_{8}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{5,7}$ (Line 4): As proven in (Maasch et al. 2024), the independence relations tested in Line 4 define partition $\mathbf{Z}_{7}$ in all structures and $\mathbf{Z}_{5}$ in structures for which $|\mathbf{Z}_{1}|=0$ (i.e., no confounders, latent nor observed, exist for $X$ and $Y$). Therefore, this test is asymptotically guaranteed to return all $\mathbf{Z}_{7}$ for arbitrary structures, and all $\mathbf{Z}_{5}$ if $\mathbf{Z}_{1}$ is the empty set.
  • $\mathbf{Z}_{4}$ (Line 5): As proven in (Maasch et al. 2024), the independence relations tested in Line 5 uniquely define partition $\mathbf{Z}_{4}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ (Line 8): At this stage, the only partitions remaining under consideration are $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{5}$ (assuming that $\mathbf{Z}_{1}$ is not the empty set). By definition, no $\mathbf{Z}_{5}$ will ever be adjacent to $Y$, as this partition cannot share a causal path to $Y$ that is not mediated by $X$ (Hernán and Robins 2006; Lousdal 2018; Maasch et al. 2024). All $\mathbf{Z}_{2\notin de(Y)}$ are descended from $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and/or $\mathbf{Z}_{4}$ (i.e., any partition that can enter into $Y$), but not from $Y$ itself. Thus, all $\mathbf{Z}_{2}$ and $\mathbf{Z}_{5}$ will be eliminated by eliminating variables not adjacent to $Y$ given $X\cup\mathbf{Z}_{4}\cup\{\mathbf{Z}^{\prime}\setminus Z\}$. This test will also result in the elimination of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are not adjacent to $Y$, including proxy confounders and proxy mediators as defined in Maasch et al. (2024).
  • $\mathbf{Z}_{4\in pa(Y)}$ (Line 10): Any intervening node on the causal path from $\mathbf{Z}_{4}$ to $Y$ ($\mathbf{Z}_{4}\to\cdots\to Y$) must also belong to a partition that can enter into $Y$ ($\cdots\to Y$). By definition, this is strictly $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, to obtain $d$-separation between $Y$ and any member of $\mathbf{Z}_{4}$ that is not adjacent to $Y$, a correct conditioning set will contain $X$ and members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$ and $\mathbf{Z}_{4}$. Conditioning on $X$ is essential, as the remaining conditioning set can open noncausal paths between $\mathbf{Z}_{4}$ and $Y$. It is therefore sufficient to condition on all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are already known to be adjacent to $Y$, unioned with $X$ and all known $\mathbf{Z}_{4}$. Any variable that is marginally dependent on $Y$ given this conditioning set is a member of $\mathbf{Z}_{4}$ that is adjacent to $Y$.
@@
  • SDC (Line 12): By definition of the SDC (Definition 3), $SDC=\mathbf{1}(X\in pa(Y))$. If $X$ is not a parent of $Y$, it will always be $d$-separable from $Y$ when all their confounding and mediating paths are blocked. If $X$ is a parent of $Y$, $X$ and $Y$ will be always conditionally dependent given any conditioning set. Therefore, the SDC will evaluate to 0 if and only if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$.
B.2 Proof of Theorem 2
@@
Maasch et al. (2024) have shown previously that the marginal and conditional independence tests used to identify $\mathbf{Z}_{4}$ and $\mathbf{Z}_{8}$ are not affected by latent variables. Intuition for this follows from the fact that these tests only require knowledge of $\{X,Y\}$ and the candidate variable $Z$. In Algorithm 1, the test for $\mathbf{Z}_{5,7}$ also relies only on knowledge of $X$, $Y$, and $Z$, and is therefore correct in the presence of latent variables. As we do not require identification of $\mathbf{Z}_{5}$ itself for downstream results, this test suffices. Otherwise, latent variables could pose identifiability problems for $\mathbf{Z}_{5}$, as in prior works (Maasch et al. 2024). A2 alone is sufficient to ensure $d$-separation between $Y$ and its non-parents at Lines 8 and 10, guaranteeing identification of the true parents of $Y$ in $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, latent variables that are not parents of $Y$ have no impact on Algorithm 1. ∎
B.3 Proof of Theorem 3
@@
This theorem claims that $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ is a VAS for WCDE identification under the conditions described in Section 2.2. Per the identifiability conditions of the WCDE (Definition 5), a VAS must block all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$. As shown in Pearl (2001), knowledge of all mediators for $\{X,Y\}$ that are parents of $Y$ is sufficient for identifying the CDE (i.e., knowledge of every mediator in the full causal graph is unnecessary). Thus, we show that adjusting for $\mathbf{A}_{\text{{DE}}}$ (1) blocks all backdoor paths for $\{X,Y\}$, (2) blocks all backdoor paths for $\{\mathbf{M},Y\}$, (3) blocks all frontdoor paths for $\{X,Y\}$, and (4) does not open non-causal paths by conditioning on colliders. We use the language of causal partitions defined in Maasch et al. (2024).
@@
    Backdoor paths for $\{X,Y\}$ are blocked. Any confounder for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{1}$, by definition of $\mathbf{Z}_{1}$ (Table 1; Maasch et al. 2024). All backdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all confounders that are parents of $Y$, all backdoor paths for $\{X,Y\}$ will be blocked when conditioning on $\mathbf{A}_{\text{{DE}}}$.
@@
    Backdoor paths for $\{\mathbf{M},Y\}$ are blocked. Any confounder for $\mathbf{M}$ and $Y$ will be in $\mathbf{Z}_{1}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{4}$ (e.g., Figure B.1), as these are the only causal partitions with edges pointing into $Y$. Set $\mathbf{A}_{\text{{DE}}}$ by definition contains all members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are parents of $Y$.
@@
    Frontdoor paths for $\{X,Y\}$ are blocked. Any mediator for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{3}$, by definition of $\mathbf{Z}_{3}$ (Table 1; Maasch et al. 2024). All frontdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all mediators that are parents of $Y$, all frontdoor paths will be blocked.
@@
    Non-causal paths through colliders are blocked. Though $\mathbf{Z}$ can contain colliders in $\mathbf{Z}_{2\notin de(Y)}$, these cannot be parents of $Y$ by definition. As defined, $\mathbf{A}_{\text{{DE}}}$ does not contain any members of $\mathbf{Z}_{2}$. Thus, non-causal paths will not be opened by conditioning on $\mathbf{A}_{\text{{DE}}}$.
@@
[…]
Figure B.1: $Y$ and members of $\mathbf{Z}_{3}$ can be confounded by members of $\mathbf{Z}_{1}$ (e.g., left), $\mathbf{Z}_{3}$ (e.g., center), or $\mathbf{Z}_{4}$ (e.g., right), as these are the only causal partitions with edges pointing into $Y$.
Appendix C Beyond Bias: Variance and Actionability
Definition 8 defines a VAS for the WCDE (Equation 5) in the general setting, irrespective of A1 and A2. However, in settings that assume A1 and A2, bypassing structure learning and adjusting for all $\mathbf{Z}$ does not induce bias in WCDE estimation. However, this is not advised for two reasons: (1) this risks unnecessary adjustment, which can inflate the variance of causal effect estimators under finite data (Schisterman, Cole, and Platt 2009) and (2) ignoring causal structure limits the informativeness and actionability of fairness conclusions.
Impacts on Asymptotic Variance
While $\mathbf{Z}$ itself constitutes a VAS under Assumptions A1 and A2, adjusting for all $\mathbf{Z}$ with no knowledge of its causal structure can undermine statistical efficiency. We highlight some prior theorems here, using the language of causal partitions.
Consider the optimality of a VAS with respect to the asymptotic variance (AV) of the causal effect estimator. For total effect estimation in both linear and nonparametric settings, it has been proven that the asymptotically optimal VAS captures less or equal information about the exposure and more or equal information about the outcome than alternative VAS (Rotnitzky and Smucler 2020; Witte et al. 2020; Henckel, Perković, and Maathuis 2022). This implies: [^13] See Theorem 1 and its corollaries in Henckel, Perković, and Maathuis (2022), which were later proven for the nonparametric case in Rotnitzky and Smucler (2020).
@@ Appendix C Beyond Bias: Variance and Actionability
    VAS containing members of $\mathbf{Z}_{5}$ will have greater or equal AV than VAS not containing any $\mathbf{Z}_{5}$;
@@
    VAS containing members of $\mathbf{Z}_{4}$ will have lesser or equal AV than VAS not containing any $\mathbf{Z}_{4}$; and
@@
    VAS containing $\mathbf{Z}_{1\not\in pa(Y)}$ or $\mathbf{Z}_{4\not\in pa(Y)}$ will have greater or equal AV than VAS where all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{4}$ are in $pa(Y)$.
While these results apply to total effect estimation, we follow this intuition for WCDE estimation using $\mathbf{A}_{\text{{DE}}}$, which excludes $\mathbf{Z}_{5}$ and $\mathbf{Z}_{1,4\not\in pa(Y)}$. We defer formal derivation of the asymptotically optimal VAS for the WCDE to future work. In Section 5 and Appendix D.5, we empirically demonstrate variance inflation in finite samples when adjusting for all $\mathbf{Z}$ versus $\mathbf{A}_{\text{{DE}}}$.
Actionable Fairness Conclusions
Ultimately, CFA must support our ability to take actions that remove unwanted association between the protected attribute and outcome. In CFA, we examine causal structure to (1) determine whether unfairness is present and (2) identify structural mechanisms that can be redressed through interventions (e.g., policy change). LD3 addresses goal (1) by using the SDC and WCDE for binary classification of the SCM (fair/unfair) with respect to direct discrimination. Through causal partition labeling, LD3 yields interpretable VAS that provide partial information for goal (2). Adjusting for all $\mathbf{Z}$ in WCDE estimation with no prior structure learning can address (1), but not (2). Consider the following example SCMs with unknown graphical structure, where A1 and A2 hold.
  1. 1.
    (A) WCDE $\neq$ 0, (B) $\mathbf{Z}_{1,3}$ is empty. We can assess (A) by adjusting for $\mathbf{Z}$ without structure learning, but we cannot know (B). Thus, we can conclude that the SCM is unfair with respect to direct discrimination, but not that indirect and spurious unfairness are absent. Discovery methods like LD3 address (A) and (B) by enabling precise WCDE estimation and indicating that policy interventions need only address direct mechanisms of unfairness.
  2. 2.
    (A) WCDE = 0, (B) $\mathbf{Z}_{1,3}$ is not empty. All pathways of unfairness in the SCM pass through $\mathbf{Z}_{1,3\in pa(Y)}$. We can ask the policy design question: are any variables in $\mathbf{Z}_{1,3\in pa(Y)}$ feasible points of intervention for mitigating unfairness? While (A) implies (B) even without structure learning, we cannot answer our policy design question without the structural knowledge returned by methods like LD3.
Empirical demonstration of improved VAS interpretability is given in Section 5. Examples of structural insights provided by local discovery are given in the real-world CFA described in Section 6.
Appendix D Empirical Validation on Synthetic and Semi-Synthetic Data
Computing Resources
All experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
D.1 LD3 in the Presence of Latent Variables
We explored the impacts of latent variables that do and do not violate Assumption A2 (Table D.1). We iteratively dropped individual variables from the observed data represented by Figure D.1, including variables that do and do not act as confounders.
Results illustrate that (1) Theorem 2 holds and (2) Assumption A2 is sufficient but not necessary. WCDE for latent confounders $B_{1}$, $M_{1}$, and $Z_{4a}$ remain unbiased, as these are not adjacent to $Y$ (Theorem 2). Though latent $M_{2}$ could induce bias by allowing for the erroneous inclusion of $M_{3}$ in $\mathbf{A}_{\text{{DE}}}$, WCDE results are not significantly impacted in this setting. Additionally, though $Z_{3c}$ is a parent of $Y$, dropping this variable simply causes LD3 to place its parent $Z_{3b}$ in $\mathbf{A}_{\text{{DE}}}$. In this setting, the predicted $\mathbf{A}_{\text{{DE}}}$ has 100% F1, precision, and recall with respect to the expected VAS, and the WCDE remains unbiased ($1.25[1.24,1.26]$). Thus, Assumption A2 is sufficient but not necessary for returning a VAS.
Latent WCDE SDC Acc $\mathbf{A}_{\text{{DE}}}$ F1 $\mathbf{A}_{\text{{DE}}}$ Prec $\mathbf{A}_{\text{{DE}}}$ Rec
$Z_{1}\in\mathbf{Z}_{1}$ 1.70 [1.66,1.75] 1.0 0.80 [0.80,0.80] 0.67 [0.67,0.67] 1.00 [1.00,1.00]
$B_{1}\in\mathbf{Z}_{1}$ 1.25 [1.24,1.26] 1.0 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00]
$B_{2}\in\mathbf{Z}_{1}$ 1.22 [1.17,1.27] 1.0 0.91 [0.89,0.93] 0.84 [0.81,0.86] 1.00 [1.00,1.00]
$B_{3}\in\mathbf{Z}_{1}$ 1.72 [1.68,1.76] 1.0 0.82 [0.8,0.840] 0.70 [0.67,0.73] 1.00 [1.00,1.00]
$M_{1}\in\mathbf{Z}_{5}$ 1.25 [1.24,1.26] 1.0 0.99 [0.98,1.01] 0.99 [0.96,1.01] 1.00 [1.00,1.00]
$M_{2}\in\mathbf{Z}_{4}$ 1.26 [1.25,1.26] 1.0 0.86 [0.86,0.86] 0.75 [0.75,0.75] 1.00 [1.00,1.00]
$Z_{4a}\in\mathbf{Z}_{4}$ 1.25 [1.24,1.26] 1.0 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00]
Table D.1: Results with latent confounders. WCDE estimates (WCDE), SDC accuracy (Acc), $\mathbf{A}_{\text{{DE}}}$ F1, $\mathbf{A}_{\text{{DE}}}$ precision (Prec), and $\mathbf{A}_{\text{{DE}}}$ recall (Rec) for LD3 as confounders in Figure D.1 are iteratively dropped from the observed data (linear-Gaussian data; Fisher-z tests; $\alpha=0.01$; true direct effect = 1.25). Values are means over 10 replicates with 95% confidence intervals in brackets. Latent nodes in red are adjacent to outcome $Y$, violating Assumption A2. $\mathbf{A}_{\text{{DE}}}$ metrics were scored by comparing the set returned by LD3 to the true $\mathbf{A}_{\text{{DE}}}$ set-minus the latent variable.
[…]
Figure D.1: Eighteen-node DAG used to evaluate the ability of LD3 to recover the true $\mathbf{A}_{\text{{DE}}}$, assess the SDC, and estimate the weighted CDE. For WCDE experiments (Figures D.3, D.4), all nodes were observed. For latent variable experiments (Section D.1), one dashed node was treated as latent per iteration.
D.2 Baseline Evaluation
Baseline Evaluation for Parent Discovery
Baselines were evaluated on the basis of F1, precision, and recall with respect to the true parents of outcome $Y$ (i.e., the members of $\mathbf{A}_{\text{{DE}}}$). Only LD3, LDECC, and MB-by-MB directly return the parents of $Y$. Additionally, PC, LDECC, and MB-by-MB return results relative to the MEC. Therefore, we evaluate each baseline with at least one form of post-processing to adjust for the task of $\mathbf{A}_{\text{{DE}}}$ prediction. Post-processing is described below.
  1. 1.
    NOTEARS: Retain all parents of the $Y$ in the unique graph.
  2. 2.
    DirectLiNGAM: Retain all parents of the $Y$ in the unique graph.
  3. 3.
    PC: Retain the intersection of parent sets across DAGs in the MEC.